Fermi National Laboratory
 

Accelerator Update for February 15, 2006 - February 17, 2006

 

Wednesday February 15

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) conducting reverse proton studies, and with NuMI and MiniBooNE taking beam.

Operations began preparing the Tevatron (TeV) at 9:04 AM for a special 1x1 (one proton bunch and one antiproton bunch) low luminosity store. The proton bunch went in at 12:26 PM. The antiproton bunch went in at 12:41 PM.

Main Injector (MI) experts began a study period at 2:19 PM.

The special 1x1 store was established around 3:40 PM.

At 9:40 PM, a Booster RF station (BRF9), which had been causing problems all evening, tripped off due to a tube fault. Operators bypassed the station.

Thursday February 16

The midnight shift began with Operations monitoring the special 1x1 store and the Recycler stash, with Pbar conducting studies, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

At 1:02 AM, a MI RF station (MIRF2) tripped off for the first of four times in 1.5 hours. It reset each time.

At 10:38 AM, Operations turned off the NuMI beam due to a problem with their Horn.

At 11:01 AM, Operations turned off the Booster beam for an access to look for a vacuum leak. The access began at 12:04 PM.

Booster RF experts began working on BRF9 at 12:42 PM.

Booster experts found the vacuum leak near a vertical collimator. Vacuum technicians tightened a flange and the leak went away.

Experts completed their work on the NuMI Horn at 2:34 PM.

Recycler experts began an Electron Cooling study at 4:57 PM.

At 5:50 PM, Booster RF experts reported that they would have had to remove the Power Amplifier (PA) for BRF9 to continue troubleshooting. They decided to leave BRF9 off and came out of access.

Booster beam returned at 6:27 PM. NuMI, MTest, and MCenter resumed taking beam.

An Operator tuned the MiniBooNE beam and had it running normally by 8:18 PM.

At 8:36 PM, Operations reported that a Linac RF station (LRF5) was failing and needed a new PA. Linac experts will begin this job on Friday morning.

Friday February 17

The midnight shift began with Operations monitoring the 1x1 store and the stash, with Pbar and MI conducting studies, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

The MI studies ended at 12:30 AM.

Pbar studies ended at 5:39 AM.

At 6:15 AM, Operations turned off Linac and Booster for access and work.

The Plans for Friday and the Weekend

The plans for today are to allow techs to complete the Linac and Booster work. Tevatron Electron Lens (TEL) experts will conduct an end-of-store study. D-Zero will make an access after the TEL studies end. The plans for the rest of the day and the weekend are to resume normal stacking, stashing, and storing. The three-month shutdown begins on Monday, February 27.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2007    email Fermilab